find the angle between the vectors i+2j+3k,3i-j+2k
Answer me ASAP plssssss

Grade:11

1 Answers

Adarsh
6 years ago
Let vector A be i + 2j + 3k and vector B be 3i – j +2k 
angle between them =>  cos@ = A.B/IAI.IBI = ½
SO @ = cos-1(1/2) = 600 = pi / 6
Thanks and Regards 
Adarsh